OpenStudy (anonymous):

Pre-Calc Question: List all solutions between [ 0 , 2pi]. cos^2(x)sin(x) = sin(x)

OpenStudy (anonymous):

I thought two of the solutions would be pi/2 and 3pi/2 since I simplified it to cos^2(x) = 0

OpenStudy (anonymous):

I see a mistake I made. I subtracted sin(x) from both sides rather than dividing.

OpenStudy (anonymous):

Let me try re-working it one second

OpenStudy (anonymous):

I'm curious as to why you ended up going further with sin though. Wouldnt you just divide sin(x) from both sides and that would leave you with cos^2(x) = 1? Then go from there

OpenStudy (anonymous):

You cannot divide by sinx. You can factor out a sinx, but that's going to lead to something more.

OpenStudy (anonymous):

Move the sinx over, factor it out, and you have \[\cos ^{2}x \sin x=\sin x\] is this what you started out with?

OpenStudy (anonymous):

Yes that is what I started with

OpenStudy (anonymous):

ok, now do not divide by sinx, instead, move it over, set the entire thing to 0, and factor it...it will look like this:

OpenStudy (anonymous):

sinx(cos^2x - 1) = 0 right?

OpenStudy (anonymous):

\[\cos ^{2}x \sin x-\sin x=0\] which by facotring out a sinx becomes: \[\sin x(\cos ^{2}x-1)=0\]

OpenStudy (anonymous):

yes, now set each of those factors =0.

OpenStudy (anonymous):

sinx=0 or cos^2x-1=0

OpenStudy (anonymous):

cos^2x-1=0 factors into (cosx-1)(cosx+1)=0

OpenStudy (anonymous):

can the cos^2x be interpreted as (cosx)^2? or is it cos(x^2)?

OpenStudy (anonymous):

You can interpret cos^2x as (cosx)^2

OpenStudy (anonymous):

Thank you very much for your help!

OpenStudy (anonymous):

You are most welcome.